Solve the system by substitution. Enter you answer as an ordered pair, or type none or infinite.
y=2x+3
y=4x-1

Answer by rothauserc(4718)

y = 2x + 3
y = 4x - 1
solve first equation for x
x = (y - 3) / 2
substitute for x in second equation
y = 4 * ((y - 3) / 2) - 1
y = 2y - 6 - 1
y = 7
substitute y = 7 in first equation
7 = 2x + 3
2x = 4
x = 2
solution is (2, 7)
**************************************************************
check solution by substituting 2 for x and 7 for y in second equation
7 = 4(2) - 1
7 = 7
solution checks